# Reconnaissance with JWST of the J-region Asymptotic Giant Branch in Distance Ladder Galaxies: From Irregular Luminosity Functions to Approximation of the Hubble Constant

### 09 Jan 2024 v1

@@ 3 JAGB Reconnaissance: Tests within Hosts
We first define a reference JAGB LF in the outer disk of each host before proceeding to divide the measures within a host for testing. For each host, we select stars with the same color cuts described in Section 2 (1 to 1.5 mag) and outside ellipses which contain the main disks of the host (see Figure 3) with parameters and magnitude cuts listed in Table 7. We discuss these choices of ellipse parameters and the effect of the measured JAGB with radius in Appendix B. We plot the CMDs for all host galaxies in Fig. 5 together with the luminosity function in the J-region and its mode-based magnitude; we also list the differences between mode and mean, the skewness, and the number of stars in that region.
We can test the stability of the JAGB by comparing fields around each host that are relatively equidistant from their main disks. The observation orientations for each host (see Figure 3) provide for such a test by comparing the fields 180 degrees apart for NGC 4258, NGC 5643 and NGC 1448. For NGC 5584 and NGC 1559 we will compare fields around the circumference of each disk.
@@
| Both sides 5% narrower | Both sides 5% wider | Blue side 5% redder | Red side 5% bluer |
@@ 3.1 NGC 4258
The CMDs for the Inner case can be found in Fig. 4, and we plot the CMDs for the Outer and Inner+Outer cases in Fig. 6. From these figures, we can immediately see an asymmetry (skew 0.4, similar to the LMC) in the J-region luminosity functions at a level which produces a difference of 0.3 mag between the mode and mean, as shown in the upper right hand corners of each subplot. We also plot the differences between the measured JAGBs in the East and West fields for all three cases, together with measurement and selection variants, in Fig. 7. For the Outer fields model fits only, the sparsity of stars makes it difficult to achieve convergence and so we bin the luminosity function with widths of 0.2 mag instead of widths of 0.1 mag. From Fig. 7, we find that the Outer fields show variations between the sides of NGC 4258 on the order of 0.1-0.2 mag for all methods; these are larger than their statistical uncertainties, indicating that the differences are intrinsic. The Inner fields show better agreement, although agreement to within their statistical uncertainties depends on measurement and selection choices. It was noted by Lee et al. (2023) that the JAGB magnitude may converge at greater radii from the host, but this may run counter to the sense here of the better agreement for the Inner fields. In the interest of greater field-to-field averaging, we will use the combined Inner+Outer fields from both sides to calibrate other host JAGB measurements (see Section 4 for more discussion).
@@ 3.2 NGC 1448 & 5643
Next, we repeat the symmetry null test described in Section 3.1 for NGC 1448 and NGC 5643. For NGC 1448, we combine photometry from visit 1, module B and Visit 2, module A, and average the magnitudes for overlapping stars. Following the recommendation of using only stars in the outer disk for JAGB measurements from Lee et al. (2022) and Lee (2023), we select stars in the outer disk that lie outside the “disk” ellipse with parameters listed in Table 7; we then further split the outer disk region into the Northwest and Southeast regions, as labeled in Fig. 3. For NGC 5643, we compare the North and South modules (visit 2 module B vs. visit 1 module A) as labelled in Fig. 3. We plot the CMDs in Fig. 8; the differences in measured JAGBs between the Northwest and Southeast fields for NGC 1448 and the North and South fields for NGC 5643 are shown in Fig. 8. We find that the region differences in the apparent JAGB reference magnitudes between the two fields within NGC 1448 and NGC 4258 are centered at approximately mag and mag, respectively, while the JAGB reference magnitudes for NGC 5643 mostly agree to within one sigma and generally 0.05 mag, suggesting that the consistency of the JAGB is better in some galaxies compared to others.
@@ 3.3 Angular Symmetry Tests
We first combine the two epochs of observations for NGC 1559 and NGC 5584 and average stars that are observed twice. We then define a window 90 wide beginning with a position angle of 0, where 0 corresponds to the unit coordinate of (1, 0). We then slide each window by 10 and repeat the JAGB measurement until lower bound of the window edges reaches 350. We repeat this test for each galaxy using four different representative measurement methods (mode with 0.25 and 0.35 mag smoothings, median, and sigma-clipped mean). We do not show the model fit as the mode with mag behaves very similarly. In addition, we choose these methods to best approximate the methodological choices from the literature (for instance, from Madore & Freedman, 2020; Ripoche et al., 2020; Lee et al., 2023). We also impose a minimum threshold of 100 stars to minimize fluctuations due to low number statistics and plot the results in Fig. 10. Yellow stars mark windows that do not overlap and are uncorrelated. Statistical uncertainties are calculated using the error on the mean. We observe differences of 0.1 mag that are not consistent with statistical uncertainties and persist across measurement methods, again suggesting some intrinsic differences in the measured JAGB.
@@ 4 Distance Ladder and
In this section, we measure distances to NGC 1448, NGC 1559, NGC 5584, and NGC 5643 using the JAGB methods defined in the previous sections. For each measurement method, we compare the SN host JAGB magnitudes to those from the combined NGC 4258 Inner+Outer fields, which best approximates the outer disk, and use the geometric maser distance of = 29.397 0.032 mag from Reid et al. (2019) to calibrate the SN host distances. The JAGB magnitudes for NGC 4258 across all variants are listed in Table 8 and plotted in the leftmost panel of Fig. 11. The other four panels of Fig. 11 give the corresponding distance modulus to each of the four galaxies for all variants, with the average shown as a red horizontal line.
For reference, we also give the distance moduli obtained from Cepheids (blue), TRGB (yellow), and Miras (cyan) where available; the Cepheid distances are obtained from the solution in (Riess et al., 2022). To eliminate calibration differences we adjust these distances upwards by 0.02 mag to reflect the distance scale based on NGC 4258 as the sole anchor. The resulting Cepheid distance moduli are = 31.31 mag, = 31.51 mag, = 31.79 mag, and = 30.56 mag, with the solid blue lines. The Mira distance to NGC 1559 is mag (Huang et al., 2020). The TRGB distances of = 31.38 mag and = 30.47 mag are from the Extragalactic Distance Database (EDD) (Tully et al., 2009; Anand et al., 2021), while = 31.49 mag and = 31.80 mag are from JWST measurements from Anand et al. (ApJ, submitted).
@@
| 22.539 | 0.027 | Median | Narrow | nan |
| 22.594 | 0.027 | 3 Clipped Mean | Narrow | nan |
| 22.372 | 0.027 | Mode | Narrow | 0.25 |
| 22.432 | 0.027 | Mode | Narrow | 0.35 |
| 22.419 | 0.044 | Model Fit | Narrow | nan |
| 22.568 | 0.024 | Median | Baseline | nan |
| 22.611 | 0.024 | Clipped Mean | Baseline | nan |
| 22.392 | 0.024 | Mode | Baseline | 0.25 |
| 22.452 | 0.024 | Mode | Baseline | 0.35 |
| 22.445 | 0.042 | Model Fit | Baseline | nan |
| 22.606 | 0.023 | Median | Wide | nan |
| 22.641 | 0.023 | Clipped Mean | Wide | nan |
| 22.441 | 0.023 | Mode | Wide | 0.25 |
| 22.511 | 0.023 | Mode | Wide | 0.35 |
| 22.516 | 0.034 | Model Fit | Wide | nan |
| 22.539 | 0.027 | Median | Blue 5 | nan |
| 22.595 | 0.027 | Clipped Mean | Blue 5 | nan |
| 22.372 | 0.027 | Mode | Blue 5 | 0.25 |
| 22.432 | 0.027 | Mode | Blue 5 | 0.35 |
| 22.424 | 0.042 | Model Fit | Blue 5 | nan |
| 22.568 | 0.025 | Median | Red 5 | nan |
| 22.61 | 0.025 | Clipped Mean | Red 5 | nan |
| 22.382 | 0.025 | Mode | Red 5 | 0.25 |
| 22.452 | 0.025 | Mode | Red 5 | 0.35 |
| 22.442 | 0.044 | Model Fit | Red 5 | nan |
@@
Next, we average over the four hosts and compare the median JAGB distances and Cepheid-based distances as shown in Fig. 12 for each measurement variant. This sample difference allows us to determine the value of relative to the complete distance ladder provided in Riess et al. (2022), i.e., allowing for this single, measured offset between this sample of differences between the first and second rungs, .
@@
Unfortunately, as seen in Fig. 12, this sample difference does depend quite significantly on which measurement method we choose, with a full range of 0.18 mag. In the next section we identify the source of this methodological uncertainty as originating from the varying degree of asymmetry of the JAGB LFs. All of the luminosity functions exhibit to varying degrees a positive skew and negative difference between the mode and mean-based JAGB measurements, as shown in Fig. 5. The fact that the calibrator, NGC 4258, has the largest skew (but not unusual in degree as it is comparable to that of the LMC and NGC 6822), with a mean minus mode of 0.29 mag (versus the mean SN host with mean minus mode of 0.1 mag), produces a systematic bias that depends on the measure employed.
@@
Despite the method-dependent bias, the JAGB observations in the anchor NGC 4258 and four SN Ia host galaxies still provide interesting constraints on the Hubble constant. We may adopt a fiducial result and uncertainty from the median of the differences between the JAGB and Cepheid distances as shown in Fig. 12 to derive relative to that in Riess et al. (2022), i.e. 5log( / ) = (Cepheids - JAGB). Calculating relative to that from Riess et al. (2022) allows us to establish a measurement using the same treatment of SN Ia from Riess et al. (2022) and isolate differences in due to only differences in JAGB and Cepheid distances between rung one and two. We find (stat) 2.3 (sys) ( 3.2 combined) km s Mpc with a range from 72 78 km s Mpc from variations in the choice of measurement method. We list the error budget in Table 9. In the Discussion we consider whether there is compelling evidence to select any specific method over the spread of methods as to indicate a better result.
@@
| JAGB Mean four host Measurement | 0.04 | |
| Cepheid Mean four host Measurement | 0.04 | |
| Rest of DL: SN Ia, local and Hubble Flow | 0.025 | |
@@ 5 Discussion
Our findings are supported by the results of Parada et al. (2021), who had also found both variations in the JAGB luminosity and differences in the shape of J-region luminosity functions via skew, with the LMC and NGC 6822 being more asymmetric than the SMC or IC 1613. Likewise, Lee et al. (2022) identified variations in the asymmetry and luminosity of the JAGB LF in M33 which varied with radial distance from the host with the LF being most symmetric in a “sweet spot” (region 3, between regions 1, 2 and 4). In Appendix B, we present further study of the radial differentiation of the JAGB.
In Fig. 13 we compare the skewness of the JAGB luminosity functions examined in this paper by fitting them with the Lorentzian model described in Parada et al. (2021). We also include the LF for the LMC and SMC with H-band photometry retrieved from Macri et al. (2015) and Ripoche et al. (2020), respectively. The legend provides the Fisher-Pearson skew and the difference between the JAGB measured using the mode of the fit and the 3 clipped mean. We offset the LFs in magnitude based on their modes (left panel) or their sigma-clipped mean (right panel); the difference clearly illustrates how mode- and mean-based measurements differ for skewed LFs. We note that the band JAGB can also be skewed by similar amounts, as shown in Lee (2023); the same holds for the J-band LF of the LMC (Parada et al., 2021). This shows that the skew effect is not unique to a particular NIR band. In this comparison, we use the Lorentzian fit to the LFs, because GLOESS smoothing tends to naturally reduce the LF skew; this can be seen in the Appendix A. As a result, a larger smoothing window used before calculating the mode produces a result closer to the mean and median measurements (see for instance, Fig. 12). The source of variation in inferred primarily comes from the difference in the skews between NGC 4258 and host galaxies. In NGC 4258, the difference between the mode and mean is 0.29 mag, while in the host galaxies, it is 0.1 mag resulting in a 0.2 mag difference across the distance ladder.
@@ 5.1 Asymmetry in the Clouds
It is instructive to further scrutinize the JAGB LFs in the Magellanic Clouds for further guidance, as these exhibit a disparity in asymmetry similar to our target galaxies, while having the virtue that their relative distances are well-established geometrically and they have been well studied in the literature using various measurement methods. In Table 10 we provide JAGB reference magnitudes in the LMC and SMC taken from the literature, together with our measurements using the same suite of methods used in our analysis (i.e., mode with , sigma-clipped mean and the median) on the same datasets as prior authors. For the LMC, which has an asymmetric JAGB LF (Parada et al. (2021) reports a skew of ), past studies of the J vs CMD from several authors (Madore & Freedman, 2020; Freedman & Madore, 2020; Ripoche et al., 2020; Parada et al., 2021; Zgirski et al., 2021) give an apparent reference magnitude of , a range of 0.15 mag; the difference between these results is much larger than the typical quoted statistical uncertainty of 0.004 mag, which is derived from the standard deviation of the distribution (0.2–0.3 mag) divided by the square root of the number of stars (typically several thousands). With our measurements, we similarly find that the asymmetry produces a large, method-dependent range of 0.1 mag. The choice of smoothing for the mode-based method also slightly shifts the JAGB reference magnitude. Another example of this effect can be seen in Appendix A. The SMC, which has a symmetric luminosity function, has a far more stable JAGB reference magnitude that also does not change with different smoothing values. We find that the mean and median for the J-region luminosity function using the Macri et al. (2015) differ by 0.06 mag.
@@
We can check which method provides a more standardized reference magnitude by comparing the differences between the geometric distances and JAGB reference magnitudes between the SMC and LMC from Graczyk et al. (2020) and Pietrzyński et al. (2019), restrict the SMC sample to stars within 1 degree within the SMC center to cover the same region as the DEBs while avoiding the complexity of the depth of the SMC outside its core region. We measure the JAGB reference magnitudes using the same variants described above and plot the differences between the SMC and LMC reference magnitudes in Fig. 14. We shade in gray the area corresponding to the difference between the SMC and LMC geometric distances of 0.500 0.017 mag. We find that in general, the median and clipped mean methods agree most closely with the geometric differences; these methods also tend to produce the least variation in the null and angular tests shown in Section 3.3. Methods which try to measure the peak of the LF, such as the mode (especially with little LF smoothing) or a model fit to the peak, are in significantly larger disagreement with the geometric difference.
@@
Non-uniformity in the color-dependence of the luminosity function can also introduce some variation in the standard candle. In all JWST host galaxies we found that the reference magnitude varies somewhat with color, not linearly, but tending to be brighter near the middle of the color range. If the color dependence is uniform across hosts, it will have no impact on distance measurements determined by comparing different hosts. In practice, the relation between the reference magnitude and color appears less important than the shape of the luminosity function based on our comparisons of different color ranges.
@@
| LMC | Z21 | K09 | Model Fit | 12.38 | 0.11 | G20 | 18.477 | P19 | -6.21 |
| LMC | R20 | R20 | ML Median | 12.31 | 0.11 | G20 | 18.477 | P19 | -6.28 |
| LMC | P21 | R20 | Mode (Lorentzian model) | 12.23 | 0.11 | G20 | 18.477 | P19 | -6.36 |
@@
| SMC | MF20 | S06 | Selected range mean | 12.81 | 0.03 | NED | 18.965 | G20 | -6.18 |
| SMC | Z21 | K09 | Model Fit | 12.85 | 0.07 | G20 | 18.977 | S16 | -6.20 |
| SMC | R20 | R20 | ML Median | 12.87 | 0.07 | G20 | 18.96 | S16 | -6.16 |
| SMC | P21 | R20 | Mode (Lorentzian model) | 12.90 | 0.07 | G20 | 18.96 | S16 | -6.15 |
| SMC | Here | R20 | Median | 12.83 | 0.04 | S21 | 18.977 | G20 | -6.19 |
| SMC | Here | R20 | Mode = 0.25 | 12.85 | 0.04 | S21 | 18.977 | G20 | -6.17 |
| SMC | Here | R20 | Mode = 0.35 | 12.85 | 0.04 | S21 | 18.977 | G20 | -6.17 |

### 29 Feb 2024 v2

@@ 3 JAGB Reconnaissance: Tests within Hosts
We first define a reference JAGB luminosity function in the outer disk of each host before proceeding to divide the measures within a host for testing. For each host, we select stars with the same color cuts described in Section 2 (1 to 1.5 mag) and outside ellipses which contain the main disks of the host (see Figure 3) with parameters and magnitude cuts listed in Table 7. We discuss these choices of ellipse parameters and the effect of the measured JAGB with radius in Appendix C. We plot the CMDs for all host galaxies in Fig. 5 together with the luminosity function in the J-region and its mode-based magnitude; we also list the differences between mode and mean, the skewness, and the number of stars in that region. We can also test for the possibility of a tilt to the J-region in the band using the Pearson correlation coefficient for the stars enclosed in each dashed box. Pearson coefficients of 1 would indicate a linear relationship, while a coefficient of 0 would indicate a lack thereof. In the top left corner of each subplot in Fig. 5, we show using the Python scipy.stats.pearsonr routine the Pearson correlation coefficient. We find that all five galaxies have Pearson correlation coefficients close to zero. (see discussion in Appendix A).
We can test the stability of the JAGB by comparing fields around each host that are relatively equidistant from their main disks. The observation orientations for each host (see Figure 3) provide for such a test by comparing the fields 180 degrees apart for NGC 4258, NGC 1448 and NGC 5643. For NGC 5584 and NGC 1559 we will compare fields around the circumference of each disk.
@@
| (All 5 used for | Both sides 5% narrower | Both sides 5% wider | Blue side 5% redder | Red side 5% bluer | |
| each method) | mag | mag | mag | mag | mag |
@@ 3.1 NGC 4258
The CMDs for these cases can be seen in Fig. 4. From these figures, we can immediately see an asymmetry (skew 0.4, similar to the LMC) in the J-region luminosity functions at a level which produces a difference of 0.3 mag between the mode and mean, as shown in the upper right hand corners of each subplot. We also plot the differences between the measured JAGBs in the East and West fields for all three cases, together with measurement and selection variants, in Fig. 6. For the Outer fields model fits only, the sparsity of stars makes it difficult to achieve convergence and so we bin the luminosity function with widths of 0.2 mag instead of widths of 0.1 mag. From Fig. 6, we find that the Outer fields show variations between the sides of NGC 4258 on the order of 0.1-0.2 mag for all methods; these are larger than their statistical uncertainties, indicating that the differences are intrinsic. The Inner fields show better agreement, although agreement to within their statistical uncertainties depends on measurement and selection choices. It was noted by Lee et al. (2024) that the JAGB magnitude may converge at greater radii from the host, but this may run counter to the sense here of the better agreement for the Inner fields. In the interest of greater field-to-field averaging, we will use the combined Inner+Outer fields from both sides to calibrate other host JAGB measurements (see Section 4 for more discussion).
@@ 3.2 NGC 1448 & 5643
Next, we repeat the symmetry null test described in Section 3.1 for NGC 1448 and NGC 5643. For NGC 1448, we combine photometry from visit 1, module B and Visit 2, module A, and average the magnitudes for overlapping stars. Following the recommendation of using only stars in the outer disk for JAGB measurements from Lee et al. (2022); Lee (2023); Lee et al. (2024), we select stars in the outer disk that lie outside the “disk” ellipse with parameters listed in Table 7; we then further split the outer disk region into the Northwest and Southeast regions, as labeled in Fig. 3. For NGC 5643, we compare the North and South modules (visit 2 module B vs. visit 1 module A) as labelled in Fig. 3. We plot the CMDs in Fig. 7; the differences in measured JAGBs between the Northwest and Southeast fields for NGC 1448 and the North and South fields for NGC 5643 are shown in Fig. 7. We find that the region differences in the apparent JAGB reference magnitudes between the two fields within NGC 1448 and NGC 4258 are centered at approximately mag and mag, respectively, while the JAGB reference magnitudes for NGC 5643 mostly agree to within one sigma and generally 0.05 mag, suggesting that the consistency of the JAGB is better in some galaxies compared to others.
@@ 3.3 Angular Symmetry Tests
We first combine the two epochs of observations for NGC 1559 and NGC 5584 and average stars that are observed twice. We then define a window 90 wide beginning with a position angle of 0, where 0 corresponds to the unit coordinate of (1, 0). We then slide each window by 10 and repeat the JAGB measurement until lower bound of the window edges reaches 350. We repeat this test for each galaxy using four different representative measurement methods (mode with 0.25 and 0.35 mag smoothings, median, and sigma-clipped mean). We do not show the model fit as the mode with mag behaves very similarly. In addition, we choose these methods to best approximate the methodological choices from the literature (for instance, from Madore & Freedman, 2020; Ripoche et al., 2020; Lee et al., 2024). We also impose a minimum threshold of 100 stars to minimize fluctuations due to low number statistics and plot the results in Fig. 9. Yellow stars mark windows that do not overlap and are uncorrelated. Statistical uncertainties are calculated using the error on the mean. We observe differences of 0.1 mag that are not consistent with statistical uncertainties and persist across measurement methods, again suggesting some intrinsic differences in the measured JAGB.
@@ 4 Distance Ladder and
In this section, we measure distances to NGC 1448, NGC 1559, NGC 5584, and NGC 5643 using the JAGB methods defined in the previous sections. For each measurement method, we compare the SN host JAGB magnitudes to those from the combined NGC 4258 Inner+Outer fields, which best approximates the outer disk, and use the geometric maser distance of = 29.397 0.032 mag from Reid et al. (2019) to calibrate the SN host distances. The JAGB magnitudes for NGC 4258 across all variants are listed in Table 8 and plotted in the leftmost panel of Fig. 10. The other four panels of Fig. 10 give the corresponding distance modulus to each of the four galaxies for all variants, with the median shown as a red horizontal line.
For reference, we also give the distance moduli obtained from Cepheids (blue), TRGB (orange), and Miras (cyan) where available; the Cepheid distances are obtained from the solution in (Riess et al., 2022). To eliminate calibration differences we adjust these distances upwards by 0.02 mag to reflect the distance scale based on NGC 4258 as the sole anchor. The resulting Cepheid distance moduli are = 31.31 mag, = 31.51 mag, = 31.79 mag, and = 30.56 mag, with the solid blue lines. The Mira distance to NGC 1559 is mag (Huang et al., 2020). The TRGB distances of = 31.38 mag and = 30.47 mag are from the Extragalactic Distance Database (EDD) (Tully et al., 2009; Anand et al., 2021), while = 31.49 mag and = 31.80 mag are from JWST measurements from Anand et al. (2024).
@@
| 22.487 | 0.018 | Median | Narrow | nan |
| 22.579 | 0.018 | 3 Clipped Mean | Narrow | nan |
| 22.35 | 0.018 | Mode | Narrow | 0.25 |
| 22.4 | 0.018 | Mode | Narrow | 0.35 |
| 22.418 | 0.019 | Model Fit | Narrow | nan |
| 22.505 | 0.017 | Median | Baseline | nan |
| 22.599 | 0.017 | 3 Clipped Mean | Baseline | nan |
| 22.36 | 0.017 | Mode | Baseline | 0.25 |
| 22.41 | 0.017 | Mode | Baseline | 0.35 |
| 22.43 | 0.019 | Model Fit | Baseline | nan |
| 22.537 | 0.015 | Median | Wide | nan |
| 22.624 | 0.015 | 3 Clipped Mean | Wide | nan |
| 22.38 | 0.015 | Mode | Wide | 0.25 |
| 22.45 | 0.015 | Mode | Wide | 0.35 |
| 22.48 | 0.019 | Model Fit | Wide | nan |
| 22.488 | 0.018 | Median | Blue 5 | nan |
| 22.582 | 0.018 | 3 Clipped Mean | Blue 5 | nan |
| 22.35 | 0.018 | Mode | Blue 5 | 0.25 |
| 22.41 | 0.018 | Mode | Blue 5 | 0.35 |
| 22.423 | 0.02 | Model Fit | Blue 5 | nan |
| 22.503 | 0.017 | Median | Red 5 | nan |
| 22.597 | 0.017 | 3 Clipped Mean | Red 5 | nan |
| 22.35 | 0.017 | Mode | Red 5 | 0.25 |
| 22.41 | 0.017 | Mode | Red 5 | 0.35 |
| 22.426 | 0.019 | Model Fit | Red 5 | nan |
@@
Next, we take the median over the four hosts and compare the median JAGB distances and Cepheid-based distances as shown in Fig. 11 for each measurement variant. This sample allows us to determine the value of .
@@
Unfortunately, as seen in Fig. 11, this sample difference does depend quite significantly on which measurement method we choose, with a full range of 0.19 mag. In the next section we identify the source of this methodological uncertainty as originating from the varying degree of asymmetry of the JAGB LFs. All of the luminosity functions exhibit to varying degrees a positive skew and negative difference between the mode and mean-based JAGB measurements, as shown in Fig. 5. The fact that the calibrator, NGC 4258, has the largest skew (but not unusual in degree as it is comparable to that of the LMC and NGC 6822 from Parada et al. (2021)), with a mean minus mode of 0.27 mag (versus the mean SN host with mean minus mode of 0.1 mag), produces a systematic bias that depends on the measure employed.
@@
Despite the method-dependent bias, the JAGB observations in the anchor NGC 4258 and four SN Ia host galaxies still provide interesting constraints on the Hubble constant. We may adopt a fiducial result and uncertainty from the median of the differences between the JAGB and Cepheid distances as shown in Fig. 11 to derive using a similar approach in Scolnic et al. (2023). We calculate the weighted mean SN Ia luminosity, , calibrated to the median JAGB distances shown in Fig. 10 and using the 6 SN Ia (SN 2001el, SN 2021pit, SN 2005df, SN 2007af, SN 2013aa, SN 2017cbv) with standardized brightnesses, , listed in Table 6 from Riess et al. (2022) to find mag. We then use this value to solve for using Equation 8 in Scolnic et al. (2023) and adopt the intercept of the Hubble diagram, from Riess et al. (2022). We find (stat) 2.3 (sys) ( 3.1 combined) km s Mpc with a range from 71 78 km s Mpc from variations in the choice of measurement method. We list the error budget in Table 9. In the Discussion we consider whether there is compelling evidence to select any specific method over the spread of methods as to indicate a better result.
@@
| JAGB Median four host Measurement | 0.04 | |
| 0.05 | ||
| Rest of DL: SN Ia, local and Hubble Flow | 0.01 | |
@@ 5 Discussion
Our findings are supported by the results of Parada et al. (2021), who had also found both variations in the JAGB luminosity and differences in the shape of J-region luminosity functions via skew, with the LMC and NGC 6822 being more asymmetric than the SMC or IC 1613. Likewise, Lee et al. (2022) identified variations in the asymmetry and luminosity of the JAGB LF in M33 which varied with radial distance from the host with the LF being most symmetric in a “sweet spot” (region 3, between regions 1, 2 and 4). In Appendix C, we present further study of the radial differentiation of the JAGB.
In Fig. 12 we compare the skewness of the JAGB luminosity functions examined in this paper by fitting them with the Lorentzian model described in Parada et al. (2021). We also include the LF for the LMC and SMC with H-band photometry retrieved from Macri et al. (2015) and Ripoche et al. (2020), respectively. The legend provides the Fisher-Pearson skew and the difference between the JAGB measured using the mode of the fit and the 3 clipped mean. We offset the LFs in magnitude based on their modes (left panel) or their sigma-clipped mean (right panel); the difference clearly illustrates how mode- and mean-based measurements differ for skewed LFs. We note that the band JAGB can also be skewed by similar amounts, as shown in Lee (2023); the same holds for the J-band LF of the LMC (Parada et al., 2021). This shows that the skew effect is not unique to a particular NIR band. In this comparison, we use the Lorentzian fit to the LFs, because GLOESS smoothing tends to naturally reduce the LF skew; this can be seen in Appendix B. As a result, a larger smoothing window used before calculating the mode produces a result closer to the mean and median measurements (see for instance, Fig. 11). The source of variation in inferred primarily comes from the difference in the skews between NGC 4258 and host galaxies. In NGC 4258, the difference between the mode (using GLOESS) and mean is 0.27 mag, while in the host galaxies, it is 0.1 mag resulting in a 0.2 mag difference across the distance ladder.
@@ 5.1 Asymmetry in the Clouds
It is instructive to further scrutinize the JAGB luminosity functions in the Magellanic Clouds for further guidance, as these exhibit a disparity in asymmetry similar to our target galaxies, while having the virtue that their relative distances are well-established geometrically and they have been well studied in the literature using various measurement methods. In Table 10 we provide JAGB reference magnitudes in the LMC and SMC taken from the literature, together with our measurements using the same suite of methods used in our analysis (i.e., mode with , sigma-clipped mean and the median) on the same datasets as prior authors. For the LMC, which has an asymmetric JAGB LF (Parada et al. (2021) reports a skew of ), past studies of the J vs CMD from several authors (Madore & Freedman, 2020; Freedman & Madore, 2020; Ripoche et al., 2020; Parada et al., 2021; Zgirski et al., 2021) give an apparent reference magnitude of , a range of 0.15 mag; the difference between these results is much larger than the typical quoted statistical uncertainty of 0.004 mag, which is derived from the standard deviation of the distribution (0.2–0.3 mag) divided by the square root of the number of stars (typically several thousands). With our measurements, we similarly find that the asymmetry produces a large, method-dependent range of 0.1 mag. The choice of smoothing for the mode-based method also slightly shifts the JAGB reference magnitude. Another example of this effect can be seen in Appendix B. The SMC, which has a symmetric luminosity function, has a far more stable JAGB reference magnitude that also does not change with different smoothing values. We find that the mean and median for the J-region luminosity function using the Macri et al. (2015) data differ by 0.06 mag.
@@
We can check which method provides a more standardized reference magnitude by comparing the differences between the geometric distances and JAGB reference magnitudes between the SMC and LMC from Graczyk et al. (2020) and Pietrzyński et al. (2019) using the LMC and SMC samples from Ripoche et al. (2020). We restrict the SMC sample to stars within 1 degree within the SMC center to cover the same region as the detached eclipsing binaries while avoiding the complexity of the depth of the SMC outside its core region. We correct for reddening using dust maps from Skowron et al. (2021) and measure the JAGB reference magnitudes using the same variants described above and plot the differences between the SMC and LMC reference magnitudes in Fig. 13. We shade in gray the area corresponding to the difference between the SMC and LMC geometric distances of 0.500 0.017 mag. We find that in general, the clipped mean method agrees most closely with the geometric differences; this method also tends to produce the least variation in the null and angular tests shown in Section 3. Methods which try to measure the peak of the LF, such as the mode (especially with little LF smoothing) or a model fit to the peak, are in significantly larger disagreement with the geometric difference.
@@
Non-uniformity in the color-dependence of the luminosity function can also introduce some variation in the standard candle. In all JWST host galaxies we found that the reference magnitude varies somewhat with color, not linearly, but tending to be brighter near the middle of the color range (see Appendix A). If the color dependence is uniform across hosts, it will have no impact on distance measurements determined by comparing different hosts as distance measurements are constructed by comparing magnitudes relative to the anchor. In practice, the relation between the reference magnitude and color appears less important than the shape of the luminosity function based on our comparisons of different color ranges.
@@
| LMC | Z21 | K09 | Model Fit | 12.38 | 0.11 | Go20 | 18.477 | P19 | -6.21 |
| LMC | R20 | R20 | ML Median | 12.31 | 0.11 | Go20 | 18.477 | P19 | -6.28 |
| LMC | P21 | R20 | Mode (Lorentzian model) | 12.23 | 0.11 | Go20 | 18.477 | P19 | -6.36 |
@@
| LMC | Here | R20 | Model Fit | 12.29 | 0.08 | S21 | 18.477 | P19 | -6.27 |
| SMC | MF20 | S06 | Selected range mean | 12.81 | 0.03 | NED | 18.965 | Gr14 | -6.18 |
| SMC | Z21 | K09 | Model Fit | 12.85 | 0.07 | Go20 | 18.977 | Gr20 | -6.20 |
| SMC | R20 | R20 | ML Median | 12.87 | 0.07 | Go20 | 18.96 | S16 | -6.16 |
| SMC | P21 | R20 | Mode (Lorentzian model) | 12.90 | 0.07 | Go20 | 18.96 | S16 | -6.15 |
| SMC | Here | R20 | Median | 12.83 | 0.04 | S21 | 18.977 | Gr20 | -6.19 |
| SMC | Here | R20 | Mode = 0.25 | 12.85 | 0.04 | S21 | 18.977 | Gr20 | -6.17 |
| SMC | Here | R20 | Mode = 0.35 | 12.85 | 0.04 | S21 | 18.977 | Gr20 | -6.17 |
| SMC | Here | R20 | Model Fit | 12.85 | 0.04 | S21 | 18.977 | Gr20 | -6.17 |
We note that Parada et al. (2021) do not use the mode from their Lorentzian fit as a final JAGB reference magnitude. -for MF20 extinction is from MW only, other studies include both MW and Cloud extinction.